OpenStudy (anonymous):

Can someone please help me with solving radical equations and expressions i need help with two problems.. one is a radical equation and the other is a radical expression i do not want you to give me the answer i need you to teach me step i really need help

OpenStudy (anonymous):

What are the expressions?

OpenStudy (anonymous):

\[(3+\sqrt{6})(3-\sqrt{6})\]

OpenStudy (anonymous):

Alright. Just distribute the terms. Your roots shouldn't be a problem on this one. Are you familiar with how to distribute?

OpenStudy (anonymous):

yes but i think dealing with the square roots throw me off and i don't exactly know what to do can you show me how to distribute a problem like this?

OpenStudy (loser66):

you have the form of (a+b)(a-b) =a^2 -b^2 , go this way, it is easier

OpenStudy (anonymous):

I'll try. Thinking about it, this is kind of tough to explain through text. So, if we had something like: (1 + 2)(3 + 4), and we weren't allowed to do the addition first, we'd distribute as normal: 1*3+1*4+2*3+2*4, and then simplify to get the obvious answer of 21. We do the same thing here, and we end up with: \[3*3+3\sqrt{6}-3\sqrt{6}-\sqrt{6}*\sqrt{6}\] Since the square root of 6 times the square root of 6 is just 6, and we can cancel the middle two terms (adding and subtracting the same thing), can simplify this to just a single number. 3 (3*3 = 9. 9 -6=3)

OpenStudy (anonymous):

Thank you it is kinda difficult to understand but i know if i just really keep practicing i will get it thank you so much for helping me and taking the time to show me an example of it

OpenStudy (anonymous):

Thank you too @loser66

OpenStudy (anonymous):

Does anyone else have another way of explaining it too all the help i cam get will be useful

OpenStudy (anonymous):

can*

OpenStudy (loser66):